EQUAL VOLUME OF TWO SOLUTIONS OF pH 2A and 12 RESPECTIVELY MIXED  , THEN RESULTANT  pH OF THE SOLUTION WILL BE WHAT?

Grade:12

1 Answers

Arun
5 years ago

let x be the vol. of both sol.

M1V1 +M2V2=Mnet Vnet

10-2 X x +10-12 X x= Mnet 2x

x(10-2 + 10-12) =Mnet 2x

10-2+10-12/2 = Mnet

Mnet

now taking -log ans will come as answer